# Rethinking Patch Dependence for Masked Autoencoders

### v1

@@ 4.4 Visualizations
To further understand the properties of the updated reconstruction objective, we devise a method to visualize the reconstruction of CrossMAE’s decoder blocks. Additionally, we visualize the weights applied to each encoder feature map to provide intuition for inter-block attention.
Visualizing Per-block Reconstruction. Rather than only visualizing the final reconstruction, we propose a more fine-grained visualization approach that allows us to precisely understand the effect and contribution of each decoder block.
Two key observations allow for such visualization. 1) Transformer blocks have residual connections from their inputs to outputs. We denote $f_{i}$ as the output of decoder $i$, $g_{i}(\cdot)$ as its residual path, with $f_{i}=f_{i-1}+g_{i}(f_{i-1})$. 2) The output of the last decoder block is processed by the reconstruction head $h$ to generate the reconstruction. Note that $h(\cdot)$ is linear because it is composed of two linear functions: a layer-norm [1] and a linear layer. Let $D$ be the decoder depth, $f_{0}$ be the input to the first decoder block and $y$ be the final reconstruction. We can recursively define $y=h(f_{D-1}+g_{D}(f_{D-1}))$. We expand it and by linearity of $h$ we have:
$\mathbf{y}$ $=h(f_{0}+g_{1}(f_{0})+\cdots+g_{D}(f_{D-1}))$
This decomposition allows us to express the reconstruction as an image stack, where the sum of all the levels gives us the final reconstruction. We present the visualization in Fig. 6 and analyze the contribution of each layer in the final reconstruction. We denormalize the output by patch mean and std. for visualization.
From Fig. 6 (a) and (b), we observe that different decoder blocks play different roles in reconstruction, with most details emerging at later decoder blocks. This supports the hypothesis for the need to get low-level information from early encoder blocks, motivating inter-block attention.
Visualizing Inter-block Attention Maps We visualize the attention maps of inter-block attention in 6(c). This shows that the CrossMAE model naturally leverages the inter-block attention to allow the later decoder blocks to focus on earlier encoder features to achieve reconstruction and allow the earlier decoder blocks to focus on later encoder features. This also motivates the need for different decoder blocks to attend to different encoder features, which is aligned with the performance gains obtained with inter-block attention.
@@ 5 Discussion and Conclusion
In this paper, we reassess the decoding mechanisms within MAE. Our exploration of MAE questions the necessity of using self-attention for reconstructing masked patches. To test this hypothesis, we design CrossMAE, an MAE-based framework that 1) uses cross-attention for reconstruction, 2) decodes a fraction of the masked patches, and 3) leverages different encoder features for reconstruction. CrossMAE shows similar performance and scaling properties as MAE while being more efficient. On one hand, CrossMAE opens the possibility to scale visual pretraining to longer contexts, especially in the setting of video pretraining, covering large swaths of in-the-wild visual data that have so far been computationally prohibitive to fully utilize. On the other hand, our investigations also give rise to worry: intuitively, self-attention among mask tokens in MAE should aid in consistent image reconstruction. However, as we show, the presence or absence of self-attention is almost uncorrelated with the quality of the learned representations in MAE. This may suggest that there exists a better way to leverage self-attention in masked visual pretraining. We hope that CrossMAE can serve as a starting point for the field to better explore the trade-off between self-attention and cross-attention for masked pretraining methods, potentially leading to truly scalable vision learners.
Acknowledgments. We thank Sophia Koepke, Yossi Gandelsman, and Qianqian Wang for their helpful discussions.
Supplementary Material
1 Implementation details
1.1 Attention Calculation
@@
To compare the attention values for mask tokens in vanilla MAE (Fig. 1), we trained a ViT-B/16 MAE for 800 epochs using the default hyperparameters provided in [31]. For each image, we randomly generate a 75% binary mask ($m$) for all tokens, with $m_{i}=1$ representing a token being masked and $m_{i}=0$ otherwise. During the forward pass of the decoder, for each self-attention operation, the attention map is stored. This means that for the default MAE, a total of 8 attention maps, each with 16 attention heads are stored. Based on the mask pattern, we calculate the outer product ($m\cdot m^{\top}$) for the self-attention among mask tokens, and $m\cdot(1-m^{\top})$ for the cross-attention from the mask token to the visible tokens. We then calculate the average across all feature maps and attention heads for self-attention and cross-attention to get the image average values. Lastly, we averaged across the entire ImageNet validation set to obtain the final values.
1.2 Inter-Block Attention
@@
We tried a few implementations for inter-block attention and found the following implementation to be the fastest and most memory-efficient. In this implementation, we combine inter-block attention for all encoder layers as a single forward pass of a linear layer. For each decoder block, we index into the output tensor to extract the corresponding feature map, and a layer norm will be applied before the feature map is fed into the decoder block. Other alternatives we tried include 1) performing separate inter-block attentions before each decoder block, and 2) 1x1 convolution on the stacked encoder feature maps.
@@
1.3 Ablation that Adds Self-Attention
In Section 4.3 (a), we propose adding self-attention back to CrossMAE as an ablation. In that particular ablation study, we analyze the effect of self-attention between the masked tokens, which can be used to improve the consistency for reconstruction. Specifically, we modify the formulation in the original transformer paper [56], where the mask/query tokens are first passed through a multi-head self-attention and a residual connection before being used in the multiheaded cross-attention with the features from the encoder. The primary difference with the vanilla transformer decoder implementation [56] is we do not perform casual masking in the multi-head self-attention. Please reference Fig. 7 for a more visual presentation of the method.
@@ 1.3 Ablation that Adds Self-Attention
Method
Mask
Ratio
Prediction
Ratio
Decoder
Depth
Interblock
Attention
[…]
MAE 0.75 0.75 8 - x - 103.45
MAE 0.75 0.75 8 - 53858 74.80
MAE 0.75 0.75 12 - 68386 93.48
CrossMAE 0.75 0.75 8 46930 69.33
CrossMAE 0.75 0.75 12 x 55358 73.75
CrossMAE 0.75 0.75 12 57987 84.05
CrossMAE 0.75 0.25 8 32055 54.88
CrossMAE 0.75 0.25 12 x 34191 52.45
CrossMAE 0.75 0.25 12 x 41630 65.80
CrossMAE 0.75 0.25 12 36805 63.23
Table 5: Pretraining runtime and GPU memory under different configurations. All time trials are conducted for 10 epochs (instead of the full 800 epochs) using 2 NVIDIA A100 GPUs. Memory is measured in MB per GPU, and runtime is measured in minutes. Note that the memory measured here contains both the encoder and decoder.
1.4 Ablation on Inter-block Attention
In Table LABEL:tab:ablation-wfm, the following cases are considered. 1 feature map (row 1) does not use inter-block attention. Each decoder block only takes the last feature map from the encoder as the keys and values. For scenarios where more than one feature map is used, the output of the patch embedding (input to the ViT) is also used.
1.5 Hyperparameters
[…]
@@ 1.5 Hyperparameters
optimizer AdamW [44]
@@
learning rate schedule cosine decay [43]
@@
optimizer momentum $\beta_{1},\beta_{2}$ = 0.9, 0.95 [12]
warm up epoch [25] 20, 40
@@
augmentation RandAug (9, 0.5) [18]
label smoothing [51] 0.1
mixup [62] 0.8
cutmix [61] 1.0
drop path [32] 0.1
@@ 1.6 Compute Infrastructure
Each of the pretraining and finetuning experiments is run on 2 or 4 NVIDIA A100 80GB GPUs. The batch size per GPU is scaled accordingly and we use gradient accumulation to avoid out-of-memory errors. ViTDet [39] experiments use a single machine equipped with 8 NVIDIA A100 (80GB) GPUs. We copy the datasets to the shared memory on the machines to accelerate dataloading. We use FlashAttention-2 [19] to accelerate attention calculation.
1.7 Runtime and GPU Memory
In this section, we provide quantitative comparisons of the effect of mask ratios, prediction ratios, and interblock attention on GPU memory usage (for both the encoder and decoder) and runtime. We provide runtimes (in minutes) of different settings of MAE and CrossMAE for 10 epochs with their corresponding GPU memory usage (in MB). All experiments here are conducted with 2 NVIDIA A100 (80GB) GPUs, with the standard hyperparameters provided above for pretraining. The results are listed in Tab. 5. Additionally, to compare against the vanilla MAE implementation, we provide rows where FlashAttention-2[19] is not enabled. Note that for MAE, the model does not fit on 2 A100 when FlashAttention-2 is not enabled, yet by adjusting the prediction ratio of CrossMAE, the model fits within the memory limit.

### v2

@@ 4.5 Visualizations
Visualizing Per-block Reconstruction. Rather than only visualizing the final reconstruction, we have two key observations that allow us to visualize the work performed by each decoder block: 1) Transformer blocks have skip connections from their inputs to outputs. 2) The final decoder block’s output goes through a linear reconstruction head to produce the reconstruction. As detailed in Appendix D, we can factor out each block’s contribution in the final reconstruction with linearity.
This decomposition allows expressing the reconstruction as an image stack, where summing up all the levels gives us the final reconstruction. As shown in Figure 5 (a,b), we observe that different decoder blocks play different roles in reconstruction, with most details emerging at later decoder blocks. This justifies the need for low-level features from early encoder blocks, motivating inter-block attention.
Visualizing Inter-block Attention Maps. As shown in the visualizations of the attention maps of inter-block attention in 5(c), CrossMAE naturally leverages the inter-block attention to allow the later decoder blocks to focus on earlier encoder features to achieve reconstruction and allow the earlier decoder blocks to focus on later encoder features. This underscores the necessity for different decoder blocks to attend to different encoder features, correlating with the performance improvements when inter-block attention is used.
@@ 5 Discussion and Conclusion
In our study, we present a novel understanding of MAE, demonstrating that coherent image reconstruction is achieved not through interactions between patches in the decoder but by learning a global representation within the encoder. Based on this insight, we propose replacing self-attention layers in the decoder with a simple readout function, specifically utilizing cross-attention to aggregate encoder outputs into each input token within the decoder layers independently. This approach, tested across models ranging from ViT-S to ViT-H, achieves comparable or better performance in image classification and instance segmentation with reduced computational requirements, showcasing the potential for more efficient and scalable visual pretraining methods. Our findings underscore the efficacy of the encoder’s global representation learning, paving the way for streamlined decoder architectures in future MAE implementations. CrossMAE’s efficiency and scalability demonstrate potential for large-scale visual pretraining, particularly on underutilized in-the-wild video datasets. However, our work has not yet explored scaling to models larger than ViT-H, the largest model examined in MAE, leaving this for future research.
Appendix A Implementation details
A.1 Attention Calculation
@@
To compare the attention values for mask tokens in vanilla MAE (Figure 1), we trained a ViT-B/16 MAE for 800 epochs using the default hyperparameters provided in (He et al., 2022). For each image, we randomly generate a 75% binary mask ($m$) for all tokens, with $m_{i}=1$ representing a token being masked and $m_{i}=0$ otherwise. During the forward pass of the decoder, for each self-attention operation, the attention map is stored. This means that for the default MAE, a total of 8 attention maps, each with 16 attention heads are stored. Based on the mask pattern, we calculate the outer product ($m\cdot m^{\top}$) for the self-attention among mask tokens, and $m\cdot(1-m^{\top})$ for the cross-attention from the mask token to the visible tokens. We then calculate the average across all feature maps and attention heads for self-attention and cross-attention to get the image average values. Lastly, we averaged across the entire ImageNet validation set to obtain the final values.
A.2 Attention to Masked Tokens Across Decoder Blocks
In addition to comparing the overall attention values for masked tokens in vanilla MAE, we explore whether this pattern is shared by all attention blocks. To do so, we analyze the cross-attention and self-attention magnitudes at each block of the pre-trained MAE.
In the table below, we report the attention to visible tokens and masked tokens across all decoder blocks:
Block Attention to visible tokens Attention to masked tokens
1 (closest to the encoder) 0.196 0.042
2 0.149 0.058
3 0.181 0.047
4 0.153 0.057
5 0.193 0.043
6 0.161 0.054
7 0.207 0.038
8 (closest to the reconstruction) 0.179 0.048
Sum (reported in the paper) 1.418 0.388
Table 6: Attention magnitudes for visible tokens and masked tokens to mask attention across decoder blocks.
Based on these results, we do not observe a significant increase in the attention to masked tokens as we progress through the later decoder blocks. The attention to masked tokens remains relatively consistent across all blocks. Additionally, we observe that the magnitude of attention is consistently larger for masked tokens’ cross-attention to visible tokens compared to their self-attention across all decoder layers. This pattern highlights the consistent interaction between masked tokens and visible tokens throughout the decoding process.
A.3 Inter-Block Attention
@@
We tried a few implementations for inter-block attention (IBA) and found the following implementation to be the fastest and most memory-efficient. In this implementation, we combine inter-block attention for all encoder layers as a single forward pass of a linear layer. For each decoder block, we index into the output tensor to extract the corresponding feature map, and a layer norm will be applied before the feature map is fed into the decoder block. Other alternatives we tried include 1) performing separate inter-block attentions before each decoder block, and 2) 1x1 convolution on the stacked encoder feature maps.
@@
Additionally, we further investigate the importance of using a cross-attention decoder, where each decoder block can use different feature maps from the encoder for decoding. In this experiment, we incorporated IBA into MAE, which uses only a self-attention decoder. Specifically, we concatenate the interblock attention features with the masked tokens. We then feed the combined features into MAE’s self-attention decoder. We pre-trained the model and finetuned it for Imagenet classification. The results are presented in Table. 7, where all models are pre-trained for 400 epochs. We observe that inter-block attention has negligible performance improvements for MAE, potentially because MAE only takes in one feature map in its decoder. In contrast, inter-block attention allows cross-attention layers in CrossMAE to attend to features from different encoder blocks, thanks to its decoupling of queries with keys and values.
Method Acc. (%)
MAE 83.0
MAE + IBA 83.0
CrossMAE (25%) 83.2
CrossMAE (75%) 83.3
Table 7: For MAE, inter-block attention has very small differences in terms of finetuning performance, potentially due to the fact that MAE’s decoder only takes in one set of features.
A.4 Ablation that Adds Self-Attention
In Section 4.3 (a), we propose adding self-attention back to CrossMAE as an ablation. In that particular ablation study, we analyze the effect of self-attention between the masked tokens, which can be used to improve the consistency for reconstruction. Specifically, we modify the formulation in the original transformer paper (Vaswani et al., 2017), where the mask/query tokens are first passed through a multi-head self-attention and a residual connection before being used in the multiheaded cross-attention with the features from the encoder. The primary difference with the vanilla transformer decoder implementation (Vaswani et al., 2017) is we do not perform casual masking in the multi-head self-attention. Please reference Figure 6 for a more visual presentation of the method.
@@ A.4 Ablation that Adds Self-Attention
A.5 Ablation on Inter-block Attention
In Table 3d, the following cases are considered. 1 feature map (row 1) does not use inter-block attention. Each decoder block only takes the last feature map from the encoder as the keys and values. For scenarios where more than one feature map is used, the output of the patch embedding (input to the ViT) is also used.
Additionally, we ablated an alternative to the linearly weighted inter-block attention. Instead, we enforce that the weights of the inter-block attention always sum to one. We do so by applying a softmax on the weights before performing the linear combination. We run the experiment on ViT-B with a mask ratio $p=0.75$ and a prediction ratio $\gamma=0.25$ with 400 pre-training epochs, 100 fine-tuning epochs. For this ablation, all encoder feature maps are used for inter-block attention. The results are presented in Table 8. We found this alternative design to achieve comparable results as the linearly weighted inter-block attention.
Method Acc. (%)
CrossMAE (25%) 83.2
CrossMAE (25%) w/ sftmax weights 83.2
Table 8: Softmax-weighted inter-block attention. We found that whether the weights in inter-block attention sums to one does not significantly impact the model’s performance.
In addition to the simple design of inter-block attention, we experimented with a variant of inter-block attention by further parameterizing the attention with linear projections. Specifically, rather than directly performing weighted sum aggregation to form the features for each cross-attention layer in the decoder, we added a linear projection for each encoder feature before the feature aggregation. We denote this variant as CrossMAE+LP. As shown in the Table. 9 (with ViT-B pre-trained for 400 epochs, consistent with the setting in Table. 3), adding a linear projection slightly improves the performance. This indicates that it is possible to design variants of readout functions, such as through improved inter-block attention, to improve the feature quality of CrossMAE.
Method Acc. (%)
CrossMAE 83.3
CrossMAE + LP 83.5
Table 9: Improving inter-block attention by adding linear projections to the input features. The performance gain indicates that it is possible to design variants of readout functions to improve CrossMAE.
A.6 Hyperparameters
Pre-training: The default setting is in Table 10, which is consistent with the official MAE (He et al., 2022) implementation. As mentioned in Sec. 3.4, we scale the learning rate by the ratio between mask ratio ($p$) and prediction ratio ($\gamma$) to ensure the variance of the loss is consistent with (He et al., 2022). Additionally, we use the linear learning rate scaling rule (Goyal et al., 2017b). This results in $\textit{lr}=\gamma*\textit{base\_lr}*\textit{batchsize}/(256*p)$. For Table 1, we use 12 decoder blocks, with mask ratio and prediction ratio both 75%, and interblock attention takes in all encoder feature maps. For the 400 epochs experiments in Table 2, we scale the warm-up epochs correspondingly. Other hyperparameters, such as decoder block width, are the same as MAE.
@@ A.6 Hyperparameters
optimizer AdamW (Loshchilov & Hutter, 2017)
@@
learning rate schedule cosine decay (Loshchilov & Hutter, 2016)
@@
optimizer momentum $\beta_{1},\beta_{2}$ = 0.9, 0.95 (Chen et al., 2020a)
warm up epoch (Goyal et al., 2017a) 20, 40
@@
augmentation RandAug (9, 0.5) (Cubuk et al., 2019)
label smoothing (Szegedy et al., 2016) 0.1
mixup (Zhang et al., 2018) 0.8
cutmix (Yun et al., 2019) 1.0
drop path (Huang et al., 2016) 0.1
@@ A.7 Compute Infrastructure
Each of the pretraining and finetuning experiments is run on 2 or 4 NVIDIA A100 80GB GPUs. The batch size per GPU is scaled accordingly and we use gradient accumulation to avoid out-of-memory errors. ViTDet (Li et al., 2022b) experiments use a single machine equipped with 8 NVIDIA A100 (80GB) GPUs. We copy the datasets to the shared memory on the machines to accelerate dataloading. We use FlashAttention-2 (Dao et al., 2023) to accelerate attention calculation.
Appendix B Additional Experiments
B.1 Linear Probe
We provide linear probe comparisons (at 800 epochs) for ViT-Small and ViT-Base in Table. 12. For both of these experiments, we run CrossMAE with a prediction ratio of 75% (reconstruction of all masked patches). These results show that CrossMAE achieves slightly better linear probe performance than vanilla MAE.
Method ViT-S ViT-B
MAE 49.7 65.1
CrossMAE 51.5 65.4
Table 12: Linear probe experiments of CrossMAE.
B.2 Quality of CrossMAE Reconstruction
In addition to the visual examples illustrated in Figure 2, we quantitatively evaluate the reconstruction quality of CrossMAE compared to the baseline MAE. Specifically, we measure the average Peak Signal-to-Noise Ratio (PSNR) for the reconstruction of masked patches (i.e., patches hidden from the model) on the ImageNet validation set. The results are summarized in Table 13.
Our findings indicate that CrossMAE achieves reconstruction performance closely comparable to the baseline MAE. Particularly noteworthy is that when predicting 75% of the masked patches, CrossMAE’s PSNR differs by only 0.3% from the baseline MAE. Similarly, predicting 25% of patches results in a marginal PSNR decrease of just 0.7%. These results demonstrate that CrossMAE maintains high reconstruction quality while significantly reducing computational demands.
Method Average PSNR for Masked Patches
CrossMAE (75% pred) 20.05 (-0.3%)
CrossMAE (25% pred) 19.97 (-0.7%)
MAE 20.11
Table 13: Comparing the PSNR of masked patch reconstruction on the ImageNet validation set.
B.3 Variance in Experiments
We pre-train and finetune the ViT-B with 25% prediction ratio (same setup as Table 3, i.e. pre-train for 400 epochs, finetune for 100 epochs), with four different seeds to compute a confidence interval of the result. We present the results in Table 14. The results suggest that CrossMAE has small variance between different experiments.
Seed 1 2 3 4 Avg $\pm$ std err
83.20 83.29 83.24 83.21 83.24 $\pm$ 0.018
Table 14: Results across different random seeds.
B.4 Masking Strategy
Similar to MAE (He et al., 2022), we here ablate the masking pattern. Instead of random masking, we perform grid-wise sampling that “keeps one of every four patches” (see MAE Figure 6). The finetuning performance is reported in Table. 15 for ViT-B (at 400 epochs), which shows that grid masking does not lead to additional improvements in downstream performance.
Method Acc. (%)
Grid Masking 83.2
Random Masking 83.3
Table 15: Ablation of masking strategies.
B.5 Ablations on More Downstream Datasets
To further investigate the performance of CrossMAE, we provide results on additional downstream tasks, including classification on iNaturalist2019, Places365, and semantic segmentation on ADE20K. These experiments demonstrate that CrossMAE performs comparably to MAE for transfer learning while offering improved performance on specific tasks.
We first compare the performance of CrossMAE and MAE on the iNaturalist2019 and Places365 datasets, pre-trained on ViT-B for 800 epochs.
Dataset MAE CrossMAE (0.25) CrossMAE (0.75)
iNaturalist2019 79.8 79.4 80.1
Places365 57.9 57.9 57.8
Table 16: Transfer learning results on iNaturalist2019 (Van Horn et al., 2018) and Places365 (Zhou et al., 2017), evaluated for image classification accuracy. CrossMAE performs comparably to MAE on these datasets.
To evaluate transfer learning to another task, in addition to the instance segmentation result in Table 2 on the COCO dataset (Lin et al., 2014), we also report results on ADE20K (Zhou et al., 2019) for semantic segmentation. These results further illustrate that CrossMAE learns comparable or more generalizable features compared to MAE.
Dataset MAE CrossMAE (0.25) CrossMAE (0.75)
ADE20K mIoU 47.7 47.7 48.1
Table 17: Semantic segmentation results on ADE20K. CrossMAE (0.75) shows an improvement over MAE and CrossMAE (0.25).
Appendix C Runtime and GPU Memory Comparisons with MAE
All experiments in Table 18 are conducted on a server with 4 NVIDIA A100 (80GB) GPUs, with the standard hyperparameters provided above for pretraining. NVLink is equipped across the GPUs. We use the default setting for MAE and set the global batch size to 4096. For CrossMAE, we also use the default setting with a prediction ratio 0.25, and this takes around 41GB memory per GPU without gradient accumulation (i.e., local batch size is set to 1024 samples per GPU). However, the same local batch size results in out-of-memory (OOM), which indicates that the total memory requirement is larger than the available memory for each GPU (80GB). To run MAE on same hardware, we thus employ gradient accumulation with a local batch size of 512 to maintain the global batch size. The benchmark runs each method and measures the average per epoch runtime as well as the max memory allocation for 10 training epochs. Our experiments in Figure 7 show that models with lower prediction ratios benefit more from deeper decoders. Our model performs on par or better when compared to MAE, with up to 3.7$\times$ lower decoder FLOPS.
Method
Memory
(MB/GPU)
Runtime
(min/epoch)
Acc.
MAE OOM ($>$81920) 5.19 83.3
CrossMAE 41177 3.38 83.5
Table 18: CrossMAE greatly improves the training throughput and reduces the memory requirements, lowering the barrier for masked pretraining. Statistics are measured on 4 NVIDIA A100 80GB GPUs. Please refer to Appendix C for comparison details. MAE’s default batch size exceeds the capacity of 4 GPUs, requiring gradient accumulation for runtime measurement.
Figure 7: We compare ViT-B which is pre-trained for 800 epochs with different variants of CrossMAE v.s. MAE. For CrossMAE, we vary the prediction ratio $p$ and number of decoder blocks $n$, and we denote each as ($p$, $n$). While all experiments are run with inter-block attention, CrossMAE has lower decoder FLOPS than MAE He et al. (2022) and performs on par or better.
Appendix D Visualizing the Contributions per Decoder Block
We propose a more fine-grained visualization approach that allows us to precisely understand the effect and contribution of each decoder block.
Two key observations enable per-block visualization: 1) Transformer blocks have residual connections from their inputs to outputs. Let $f_{i}$ be the output and $g_{i}(\cdot)$ the residual function of decoder $i$, so $f_{i}=f_{i-1}+g_{i}(f_{i-1})$. 2) The final decoder block’s output goes through a reconstruction head $h$, which is linear, consisting of a layer-norm and a linear layer, to produce the reconstruction. With $D$ as the decoder depth, $f_{0}$ the initial input, and $y$ the final output, $y$ is recursively defined as $y=h(f_{D-1}+g_{D}(f_{D-1}))$, which simplifies due to the linearity of $h$:
$\mathbf{y}$ $=h(f_{0}+g_{1}(f_{0})+\cdots+g_{D}(f_{D-1}))$
This decomposition allows us to express the reconstruction as an image stack, where the sum of all the levels gives us the final reconstruction. We present the visualization in Figure 5.